OpenStudy (anonymous):

Solve: 4x − 10 + 16 = 22 4 12 7 −4

OpenStudy (anonymous):

4x - 10 + 16 = 22 -10 + 16 = 6 so now you have: 4x + 6 = 22 Subtract 6 from both sides: 4x = 16 Divide both sides by 4: x = 4 Hope that helps. : )

OpenStudy (ams98):

What she said^^ :)